Determine weather is. parallel, perpendicular or neighter Yy +2=3x -3x +4y=-12
Answer #1

ao - 0 Sol Giuen lenes 44+2=34 39-44-2=0 1. Slope of the line mi =-2 / 2 = 1 44 0 <: slope of ant by te=0 i mo=97) Also ginen